### v1

# AppApp Readings

Abstract
We study doping and temperature dependencies of temporal and spatial spin stiffnesses of the Hubbard model within the mean field approach for incommensurate magnetic order. We show that the frequency dependence of temporal spin stiffness is crucial to obtain small values of correlation length, comparable to those observed in cuprates. Using the obtained spin stiffnesses, we obtain the temperature and doping dependence of correlation length within the large-$N$ limit of the respective nonlinear sigma model. In agreement with the experimental data on La2-xSrxCuO4 we obtain short range magnetic order with relatively small correlation length at $0.1\lesssim x\lesssim 0.2$, and magnetically ordered ground state in the narrow doping region $0.05\lesssim x\lesssim 0.1$. The latter state may correspond to the spin-frosen state, observed in the experimental data on La2-xSrxCuO4.
[…]
To describe the low-energy spin excitations of magnetic systems the nonlinear sigma model (NLSM) can be used, see, e.g., Ref. Auerbach, for a review. The corresponding excitations in this model are parameterized by the temporal and spatial spin stiffnesses, which can be determined from the microscopic analysis. For insulating magnets, the classical version of the $O(3)$ non-linear sigma model was derived in the continuum limit of the classical Heisenberg model, assuming commensurate antiferromagnetic order ZinnJustin ; NelsonPelkovits . Later this derivation was also generalized to the quantum case Haldane ; CHN ; CSY . The nonlinear sigma models in the $O(3)/O(2)$ (Refs. DombreRead, ; Mouhanna, ) and $CP^{N-1}$ (Refs. CPN, ; CSS, ; Azaria, ) manifolds were considered in a continuum limit of frustrated quantum antiferromagnets with spin spiral ground state.
The situation in itinerant systems is more involved. The derivation of the NLSM, which also allows obtaining the respective spin stiffnesses, was first proposed in Refs. Weng, ; Schulz, ; Dupuis, ; Dupuis1, ; Sachdev, ; Sachdev1, using Hubbard-Stratonovich transformation. This derivation introduced in the action the effective fluctuating field, corresponding to the order parameter, which is then fixed at its mean field value. More recently, using the SU(2)-symmetric gauge theory, which does not introduce symmetry breaking terms or condensation of gauge fields in the action, was proposed for derivation of the nonlinear sigma model and obtaining respective spin stiffnesses Bonetti ; BonettiThesis ; BonettiMetzner ; BonettiErratum ; OurSecond . In this approach the symmetry is broken only at the level of dressed single- and two-particle Green’s functions via spin-asymmetric self-energy and the resulting renormalized interaction vertices. The respective Ward identities for the latter approach were derived in Refs. Bonetti, ; BonettiThesis, ; BonettiErratum, ; OurSecond, . It was emphasized in Ref. OurSecond, that the temporal spin stiffness is in general dynamic, and possesses strong frequency dependence away from half filling. This frequency dependence becomes more pronounced with increase of doping.
In the present paper, we consider the doping- and temperature dependence of spatial and temporal spin stiffnesses of the non-linear sigma model, obtained within the mean field approach for incommensurate magnetic order of the Hubbard model. While previous consideration BonettiMetznerStatic ; BonettiNew was restricted to the static limit, we uncover the role of the dynamic effects in temporal spin stiffness. We emphasize that strong frequency dependence, which occurs already in the low-energy regime, corresponds to consideration of gapped magnetic modes, which are absent in the standard spin-wave approaches, but were discussed recently within the bond-operator technique Syromyatnikov . Using the obtained results, we calculate the temperature- and doping dependence of correlation length and show that the gapped modes are crucial to obtain qualitative agreement with the experimental data on high-$T_{c}$ compounds.
Model and magnetic correlations. We consider the Hubbard model on the square lattice
$H=-\sum_{i,j,\sigma}t_{ij}{c_{i\sigma}^{\dagger}c_{j\sigma}}+U\sum_{i}{n_{i\uparrow}n_{i\downarrow}},$ (1)
with hopping $t_{ij}=t$ between nearest neighbors (which is used as a unit of energy) and $t_{ij}=-t^{\prime}$ for next-nearest neighbors.
We consider spiral magnetic order in the $x,z$ plane. For the reasons described below, we consider mainly the out of plane susceptibility, which corresponds to the response $\chi^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}=\bar{\chi}^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}={-}\langle\langle S^{y}_{q}|S^{y}_{-q^{\prime}}\rangle\rangle$ (the bar denotes the local spin reference frame). This susceptibility depends on temporal and spatial spin stiffnesses $\chi_{\rm{op},\omega_{n}}$,$\rho_{\rm{op},xx/yy}$ as parameters, which can be determined from microscopic analysis, and characterizing the low energy magnetic excitations Bonetti ; BonettiErratum ; BonettiThesis ; BonettiMetzner ; OurSecond ,
$\chi^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}=\frac{m^{2}}{\Delta+\chi_{\rm{op},\omega_{n}}\omega_{n}^{2}+\rho_{\rm{op},x}q_{x}^{2}+\rho_{\rm{op},y}q_{y}^{2}},$ (2)
where $m$ is the long-range order parameter. Short-range magnetic order is characterized by spin gap $\Delta$, determining the correlation length $\xi$.
To obtain correlation length, we consider the leading order of $1/N$ expansion in the massive CPN-1 modelCPN ; CSS ; BonettiMetzner . In this case, only the out-of-plane mode of magnetic excitations contributes to the correlation length, describing low-lying magnetic excitations and short-range magnetic fluctuations, in the large number of components $N$ limit. The value of an energy gap $\Delta$ is determined in this approach by the respective sum rule CSS
$T\sum_{\omega_{n}}\iint\frac{d^{2}q}{(2\pi)^{2}}\;\chi^{yy}_{\mathbf{Q}+\mathbf{q},\omega_{n}}=\frac{m^{2}}{2}.$ (3)
where the integration is carried over sufficiently small momenta $\mathbf{q}$ (the chosen integration region is specified in Supplemental MaterialSM ), the factor 1/2 in the right-hand side accounts for the presence of two non-identical magnetic wave vectors ±$\pm{\bf Q}$ in the Brillouin zone. In the polar coordinates this sum rule results in an equation for $\Delta$
$1=T\sum\limits_{\omega_{n}}\int\limits_{-\pi}^{\pi}\frac{d\phi}{4\pi^{2}\,\rho(\phi)}\ln{\left(1+\frac{q_{m}^{2}(\phi)\rho(\phi)}{\Delta+\chi_{\rm{op},\omega_{n}}\omega_{n}^{2}}\right)},$ (4)
where $\rho(\phi)=\rho_{\rm{op},x}\cos^{2}{(\phi)}+\rho_{\rm{op},y}\sin^{2}{(\phi)}$, and $q_{m}(\phi)$ is the maximum bound of $q$-integration, see Supplemental MaterialSM . We interpret contributions from finite frequencies $\omega_{n}$ as an impact of dynamic spin fluctuations on correlation length.
Good quantitative description of an energy gap at low and intermediate temperatures can be obtained by separating the frequency sum into the possibly singular $\omega_{n}=0$ and finite frequency contributions and approximate the latter contribution, which is finite for $\Delta=0$, as an integral over frequencies. In the limit of small spin gap, we find the equation
$\Delta(T)=\bar{q}_{m}^{2}\bar{\rho}{e^{{-2\pi\kappa(T)\bar{\rho}(T)}/T}}$ (5)
with spatial spin stiffness renormalization factor $\kappa(T)$ defined as
$\kappa(T)\equiv 1-\frac{1}{\pi\bar{\rho}}\int\limits_{{\pi T}}^{+\infty}\frac{d\omega}{2\pi}\int{\frac{d\phi}{2\pi}}\ln{\left(1+\frac{q_{m}^{2}(\phi)\bar{\rho}}{\Delta(T)+\chi_{\rm{op},\omega}\omega^{2}}\right)},$ (6)
[…]
There are two distinct regimes of spin gap temperature dependence with lowering temperature. If we assume for $T\to 0$ that $\Delta\to 0$ than $\kappa(T)\to\kappa(0)\equiv\kappa>0$ and the equation for the spin gap simplifies to
$\Delta(T)\approx\bar{q}_{m}^{2}\bar{\rho}e^{-2\pi\kappa\bar{\rho}/T}.$ (7)
where
$\kappa=1-\frac{1}{\pi\bar{\rho}}\int\limits_{0}^{+\infty}\frac{d\omega}{2\pi}\int{\frac{d\phi}{2\pi}}\ln{\left(1+\frac{q^{2}_{m}(\phi)\bar{\rho}}{\chi_{{\rm op},\omega}\omega^{2}}\right)}$ (8)
In this case, the role of dynamic quantum fluctuations is only in the renormalization of spatial spin stiffness $\bar{\rho}\to\kappa\bar{\rho}$ in the exponent, similarly to the previous analysis of $O(N)/O(N-1)$, Ref. CSY, and massive $CP^{N-1}$, Ref. CSS, nonlinear sigma models. In this regime according to the equation (7) we define the crossover temperature $T^{*}=2\pi\kappa\bar{\rho}$ to the quasi-long range order state (denoted also as the renormalized classical regime CHN ), such that at $T<T^{*}$ the correlation length gets exponentially large.
Another possibility for $\Delta(T)$ is to approach a finite value with $T\to 0$. In this regime (denoted also as quantum disordered regime CHN ), the correlation length reaches a finite value with $T\to 0$ and thus the system’s ground state is disordered. Finiteness $\Delta(T)$ in the denominator of Eq. (6) is important in this regime; the equation for the zero-temperature limit of the gap $\Delta(0)$ can be obtained from Eq. (6) by putting $T=0$ and $\kappa(T=0)=0$, which implies that the renormalized spatial spin stiffness is zero in this case.
Figure 1: (a) Dynamical temporal spin stiffness for out-of-plane mode as a function of bosonic Matsubara frequency $\omega_{n}$ for various temperatures and electron concentrations. (b,c) The real frequency dependence of the imaginary part of the out-of-plane staggered dynamic susceptibility $\text{Im}\chi^{yy}_{\mathbf{Q},\omega}$ (b) and interband spectral weight $A_{{\mathbf{Q}},\rm ib}(\omega)$ (c) for different electron concentrations $n$ at temperature $T=0.025t$.
Although the above presented analysis parallels that considered previously for the NLSM (see, e.g., Refs. CHN, ; CSY, ), we emphasize the key difference, which originates from the frequency dependence of the temporal spin stiffness $\chi_{{\rm op},\omega}$. As we show below, it yields strong suppression of magnetic correlations, not allowing long range magnetic order to establish at $T=0$ and leading to the quantum disordered state. To characterize the respective tendency to short-range magnetic order, we define the crossover temperature of the onset of short-range magnetic order $T_{\rm SRO}$ which fulfills the equation $2\pi\kappa(T_{\rm SRO})\bar{\rho}(T_{\rm SRO})=T_{\rm SRO}$, such that $\xi(T_{\rm SRO})\simeq 1/\bar{q}_{m}\sim 1$. Despite similarity to the definition of the crossover temperature $T^{*}$, the temperature $T_{\rm SRO}$ has a completely different meaning and denotes the temperature below which (i.e. at $T\ll T_{\rm SRO}$) the gap becomes small, and hence the correlation length is essentially large ($\xi\gg 1$).
Mean field calculations of spin stiffnesses. For calculations, we use the standard mean-field approach; the spin stiffnesses are expressed through the spin- and current correlation functions, see Refs. Bonetti, ; BonettiMetzner, ; BonettiThesis, ; BonettiErratum, ; OurSecond, and Supplemental Material SM . We choose $t$ as a unit of energy, $t^{\prime}/t=0.16$, which is close to that for La2CuO4, and the relatively weak interaction strength $U/t=2.5$, since for stronger interaction, the incommensurate state in the mean field approach is unstable at low dopings. In view of the used mean field approach, this interaction should be, however, considered as an effective screened interaction, which is typically much smaller than its bare value, cf. Refs. KataninYamase, ; IrkhinIgoshev, . The considered interaction also coincides with that used previously in Refs. BonettiMetznerStatic, ; BonettiThesis, ; it is also close to that obtained for not too small doping range in Ref. BonettiMetzner, .
The obtained frequency dependence of the temporal spin stiffness (see Fig. 1a) is strong already for $\omega_{n}\ll 1$ (in units of $t$), where the NLSM is applicable. While the static limit varies only weakly with electron concentration $n$, the decrease of the high-frequency asymptote (which is also almost temperature-independent) corresponds to the strengthening of quantum fluctuations with lowering $n$. These quantum fluctuations are related to the finite spectral weight at finite energies. In Fig. 1b we show the resulting frequency dependence of the out-of-plane susceptibility at the real frequency axis, obtained by analytic continuation. At and close to half filling we find the peak at low frequencies corresponding to a vanishing or small spin gap. However, apart from that, we observe generally two other peaks. One can see that the position of the lower of the two peaks coincides with the low-energy edge of the inter band spectral function $A_{{\mathbf{Q}},\rm ib}(\omega)=\sum_{\mathbf{k}}\delta(E^{+}_{\mathbf{k}+\mathbf{Q}}-E^{-}_{\mathbf{k}}-\omega)(f(E^{-}_{\mathbf{k}})-f(E^{+}_{\mathbf{k}+\mathbf{Q}}))$, shown in Fig. 1, where $E_{\mathbf{k}}^{\pm}$ are the electronic energy bands. The low-energy edge of $A_{{\mathbf{Q}},\rm ib}$ is determined by an indirect gap in the electronic spectrum. This peak can be therefore identified with the Higgs mode Sachdev1 ; Sachdev2 ; Sachdev ; SachdevHiggs1 . On approaching the paramagnetic mean-field solution (i.e. at substantial doping levels) this peak shifts to very low energies, and therefore, becomes suppressed by the spin gap. Therefore, at sufficiently large dopings, we observe a single peak in the spin spectral functions. These peaks in the spin spectral function remind us also of the gapped magnetic excitations, discussed recently for the square-lattice Heisenberg model within the bond-operator technique Syromyatnikov .
The spatial spin stiffness vanishes at the electron concentration, corresponding to the transition from the antiferromagnetic to the spiral state OurSecond ; BonettiNew (see also Supplemental Material SM ). This concentration gradually shifts toward half-filling as temperature decreases yielding a rapid increase of spin stiffness near half filling, which, being continuous at a finite temperature, converges to a discontinuous function at $n=1$ in the zero-temperature limit $T\to 0$. In this limit, the spatial spin stiffness vanishes when the electron concentration approaches half-filling from the hole-doped side, while remaining finite when approached from the electron-doped side.
Figure 2: The obtained phase diagram. $T_{N}^{MF}$ (rhombs) denote the mean-field transition temperatures to the paramagnetic state, $T_{\text{inc}}$ (squares) mark the transition temperatures between spiral and antiferromagnetic (AFM) states in the mean field approach. $T_{\text{SRO}}$ (triangles) represent the temperatures at which the transition to the paramagnetic state with short range magnetic order, denoted as Spiral SRO and defined by the condition $\xi(T_{\text{SRO}})>1$, occurs in the NLSM approach. The dashed lines show contour levels of various values of correlation length (denoted by numbers at the lines in units of the lattice parameter) in the NLSM approach inside the SRO regime. $T^{*}$ corresponds to the crossover temperature to the renormalized classical (RC) regime with the exponentially large correlation length in the NLSM approach; the corresponding ground state shows a tendency to long-range order. The violet region possesses spiral order in the mean field approach, but has small correlation length $\xi\lesssim 1$ in the NLSM approach.
Phase diagram. In Fig. 2, we present the resulting phase diagram which shows crossover temperatures to the state with short- ($T_{\rm SRO}$) and quasi-long range ($T^{*}$) magnetic order, as well as regions where magnetic order is entirely destroyed by quantum fluctuations. The phase boundaries $T^{MF}_{N}$ and $T_{\text{inc}}$ represent the transitions between different long-range ordered states, as obtained within the pure mean-field framework. Near the transition between spiral and antiferromagnetic states, the long-range magnetic order is destroyed by purely static magnetic fluctuations, which arise due to the vanishing of the spatial spin stiffness, $\bar{\rho}\to 0$ (see also Ref. BonettiNew, ). This effect is absent on the electron-doping side of the phase diagram (not shown here), as the spatial spin stiffness does not vanish in that region.
Away from half-filling, the spatial spin stiffness increases, while the temporal spin stiffness at finite bosonic frequencies becomes suppressed, resulting in enhanced quantum fluctuations. At moderate doping these quantum fluctuations destroy long-range magnetic order at $T=0$, giving rise to states with short-range magnetic order characterized by a correlation length larger than the lattice parameter, $\xi>1$. These states are located in the green region of the phase diagram in Fig. 2.
At the same time, the delicate interplay of the above described effects leads to the interval of electron concentrations where the ground state exhibits long-range magnetic order that is bounded by points of vanishing of the crossover temperature $T^{*}$. The states with finite crossover temperature $T^{*}$ tend to form an ordered ground state with decreasing temperature. We emphasize in the following that the absence of long-range order at sufficiently large doping is entirely due to the account of the frequency dependence of temporal spin stiffness.
[…]
Correlation length. In order to examine the role of dynamical quantum fluctuations affecting the system via finite bosonic frequencies and dynamical temporal spin stiffness $\chi_{\rm{op},\omega_{n}}$, we consider the dependence of the magnetic energy gap, obtained from the equation (3), on electron concentration $n$. In Fig. 3 we compare results of the three different methods for $T=0.05t$: considering only the zeroth bosonic frequency $\omega_{0}$, calculation of the frequency sum with $\chi_{\omega_{n}}\approx\lim\limits_{\omega_{n}\to 0}\chi_{\omega_{n}}$, and with the dynamic $\chi_{\omega_{n}}$ approximated at high frequencies. One can see that the correlation length is decreased by almost an order of magnitude when including finite bosonic frequencies in the sum of Eq. (3). Using only the static limit of the temporal spin stiffness in this sum is, however, insufficient; suppression of the dynamic stiffnesses at high frequencies results in a further strong decrease of the correlation length. This effect of quantum fluctuations becomes strengthened with the decrease of temperature (see Supplemental Material SM ), yielding an ordered ground state when using the static limit of temporal spin stiffness. Therefore, we find that the frequency dependence of temporal spin stiffness is of crucial importance for describing spin correlations within the nonlinear sigma model.
In Fig. 3 we present also the comparison with the experimental data on doping dependence of correlation length in La2-xSrxCuO4. Choosing relatively high temperature $T=0.05t$ in theoretical result is dictated by the necessity of having AFM to incommensurate transition at substancial doping in accordance with the experimental data; is can be justified as modeling the effect of structural disorder and local correlations in the proximity of the metal-insulator transition, which are beyond the mean-field scheme. In particular, our previous study OurFirst demonstrated that strong electronic correlations can induce a redistribution of spectral weight between the lower and upper Hubbard bands. This redistribution can prevent the formation of a metallic state, thereby potentially stabilizing the antiferromagnetic ground state even at finite hole doping. For comparison, we also shift the experimental data by the doping $\Delta x=-0.025$, to adjust this transition at the same doping, as in the theoretical result. This difference of dopings may be also an effect of the presence of oxygen states. As noted in Ref. HoleDoping, , hole carriers may initially occupy nonmagnetic-energy bands rather than directly populating the antiferromagnetic band. When considering full frequency dependence of temporal spin stiffness, we find qualitative agreement with the experimental data.
In summary, we have analyzed the effect of the frequency dependence of temporal spin stiffness on magnetic order in the doped Hubbard model. We find that this frequency dependence is crucial to obtain short range magnetic order with a sufficiently small correlation length. In particular, the gapped energy modes, one of which corresponds to the Higgs mode, carry substantial spectral weight, which yields the decrease of correlation length, determined by the sum rule. As a result, we find long-range magnetic order only in the narrow domain of dopings $x\simeq 0.1$, in agreement with the experimental data. We expect that this region of long-range order may be further shrunk by fluctuations, and/or partly unstable with respect to the non-uniform patterns, see, e.g., recent study of Ref. MetznerFiniteLattice, , which finds stripe instability in approximately the same doping range.
For future studies, it would be interesting to extend the considered approach to include dynamic local fluctuations within dynamical mean field theory for incommensurate phase DMFT_Incomm_Licht ; DMFT_Incomm ; DMFT_Incomm1 ; SOurFirst ; OurSecond , which may allow considering realistic interaction strengths. Consideration of the effects of Harris disorder within the Yukawa-Sachdev-Ye-Kitaev model Harris is also of certain interest.
Note added. When finalizing the paper, we have learned about the recent study of Ref. BonettiNew, , which considered the temperature- and concentration dependencies of spin stiffnesses and correlation length. However, only static temporal spin stiffness was considered in that study.
Acknowledgements. This study is supported by the Russian Science Foundation (Grant No. 24-12-00186).
Supplemental Material
to the paper “Frequency dependence of temporal spin stiffness and short-range magnetic order
in the doped two-dimensional Hubbard model”
I. A. Goremykin and A. A. Katanin
.1 q-integration in the sumrule (3) and the choice of the cutoff parameter
Since there are generally two points ±$\pm{\mathbf{Q}}$ with softening of spin excitations, one should restrict integration in Eq. (3) to the vicinity of one of these points (the presence of two points is accounted for by the factor 1/2 in the right-hand side). The respective region of integration is shown in Fig. S1. In the radial coordinates, it implies
[…]
where $q_{m}(\phi)=\delta/\cos\phi$ for $|\phi|<\arccos(\delta/\Lambda)$ and $q_{m}(\phi)=\Lambda$ otherwise, where $\Lambda$ is a cutoff parameter. In the limit $\delta\rightarrow 0$ (i.e. close to the commensurate phase) we obtain $I=1/2\iint_{q<\Lambda}d^{2}q$, which cancels the factor 1/2 in the right-hand side of Eq. (3) and corresponds to the weight $m^{2}$ of a single mode having $q$ close to $(\pi,\pi)$.
Figure S1: Integration region (shown by colored area) in the sum rule (3)
The cutoff parameter $\Lambda$ should be chosen in the way that it cuts the short-range degrees of freedom, described by mean-field theory, keeping only the long-range ones. Let us assume for simplicity that the inverse susceptibility in the whole momentum space (bounded as described above) can be approximated as quadratic in momentum, with the respective cutoff $\Lambda_{\rm uv}$. The respective sum rule for the susceptibility reads
$\frac{\widetilde{S}^{2}}{2}=T\sum_{i\omega_{n}}\iint\limits_{0}^{\;\;\;\;\Lambda_{\rm uv}}\frac{d^{2}q}{(2\pi)^{2}}\frac{m^{2}}{\rho q^{2}+\chi\omega^{2}+\Delta}$ (A2)
where $\tilde{S}^{2}$ is the square of magnetic moment. Comparing this to the Eq. (3), neglecting $\Delta^{2}$ at $q\sim\Lambda$, and approximating frequency sum by the integral, we find
$\frac{\widetilde{S}^{2}-m^{2}}{2}=T\sum_{i\omega_{n}}\iint\limits_{\Lambda}^{\;\;\;\;\Lambda_{\text{uv}}}\frac{d^{2}q}{(2\pi)^{2}}\frac{m^{2}}{\rho q^{2}+\omega\chi^{2}}\simeq\frac{m^{2}}{4\pi\rho}\int\frac{d\phi}{2\pi}\int\frac{d\omega}{2\pi}\ln\frac{q_{\mathrm{uv}}^{2}(\phi)+(\omega/c)^{2}}{q_{m}^{2}(\phi)+(\omega/c)^{2}}$ (A3)
where $c=(\rho/\chi)^{1/2}$ and $q_{\rm uv}$ is obtained from $q_{m}$ by replacing $\Lambda\rightarrow\Lambda_{\rm uv}$. For frequency-independent temporal spin stiffness, we obtain
$\frac{\widetilde{S}^{2}-m^{2}}{2}\simeq\frac{m^{2}c}{4\pi\rho}{\int\frac{d\phi}{2\pi}}(q_{\mathrm{uv}}(\phi)-q_{m}(\phi))\simeq\frac{m^{2}}{4\pi\rho}c(\Lambda_{\mathrm{uv}}-\Lambda)$ (A4)
From this we find
$\Lambda=\Lambda_{\rm uv}-\frac{\widetilde{S}^{2}-m^{2}}{m^{2}}\frac{2\pi\rho}{c}$ (A5)
Since at $m\rightarrow 0$ we have $\rho\propto\chi\propto m^{2}$, requiring that $\Lambda$ vanishes at $m\rightarrow 0$ (which implies that all degrees of freedom are described qualitatively correct by mean-field theory in that case), we find $\Lambda_{\rm uv}=(\widetilde{S}^{2}/m^{2})\Lambda$ and $\Lambda={2\pi\rho/c}$. Interestingly, the condition $\Lambda\simeq 2\pi\rho/c$ also implies that the second term in Eq. (6) of the main text is of the order of 1, i.e. quantum fluctuations become essential for the chosen momentum cutoff. The quantity $2\pi\rho/c$ is in general temperature- and filling-dependent, but in the considered range $0.950.8<n<$ 0.95 it weakly changes with temperature and doping. Considering the average $\Lambda$, obtained with low- and high-frequency limits of temporal susceptibility, we establish the temperature- and filling-independent cutoff $\Lambda=0.6$.
.2 Mean field approach
At the mean field level for spiral spin density wave magnetic order with general magnetic wave vector ${\bf Q}$ the Hamiltonian (1) can be diagonalized and brought to the form SDW1 ; SDW2 ; SDWIc1 ; SDWIc2 ; SDWIc3 ; SDW3 ; SDW4 ; SDWOur ; OurFirst
$\hat{H}=\sum\limits_{\mathbf{k},\alpha=\pm}E_{\mathbf{k}}^{\alpha}\hat{D}_{\mathbf{k}\alpha}^{+}\hat{D}_{\mathbf{k}\alpha}+NU(m^{2}-\frac{n^{2}}{4}),$ (A6)
where the electronic energy bands are
$E^{\pm}_{\mathbf{k}}=-\tilde{\mu}+\epsilon^{+}_{\mathbf{k}}\pm\sqrt{(\epsilon_{\mathbf{k}}^{-})^{2}+\Delta^{2}_{\text{el}}},\,\,\,\epsilon^{\pm}_{\mathbf{k}}=\frac{1}{2}(\epsilon_{\mathbf{k}+\mathbf{Q}/2}\pm\epsilon_{\mathbf{k}-\mathbf{Q}/2}),$ (A7)
$\tilde{\mu}=\mu-{Un}/{2}$, and the (direct) energy gap in the electronic spectrum is $\Delta_{\text{el}}=Um+{h}/{2}$.
Figure S2: The dependence of a) local magnetization $m$ and b) normalized magnetic wave vector component $Q_{x}/\pi$ from electron concentration $n$ for different temperatures $T=\frac{1}{\beta}$. Dashed vertical lines correspond to the critical electron concentration $n_{c}$ of the transition between states with AFM and spiral long-range magnetic order.
In order to find thermodynamically stable magnetic phases, we have minimized the free energy potential with respect to the magnetic wave vector $\mathbf{Q}$. The explicit form of the free energy potential $W$ at zero external magnetic field is given by
$W=-T\ln Z=-T\sum\limits_{\mathbf{k\alpha}}\ln\left(1+e^{\frac{-E^{\alpha}_{\mathbf{k}}}{T}}\right)+NU(m^{2}-\frac{n^{2}}{4}).$ (A8)
Stable states with long-range magnetic order were found by self-consistent solution of the mean-field equations
$1=\frac{U}{2}\sum_{\mathbf{k},\alpha=\pm}\frac{-\alpha f(E^{\alpha}_{\mathbf{k}})}{\sqrt{(\epsilon_{\mathbf{k}}^{-})^{2}+\Delta^{2}_{\text{el}}}},\,\,\,n=\sum_{\mathbf{k},\alpha=\pm}f\left(E_{\mathbf{k}}^{\alpha}\right)$ (A9)
followed by the minimization of Eq. (A8) with respect to the magnetic wave vector $\mathbf{Q}$.
The computed dependence of local magnetization $m$ for states with long-range magnetic order and corresponding magnetic wave vector’s $\mathbf{Q}$ component $Q_{x}$ are presented in Fig. S2. Temperature dependence of both quantities is relatively weak, except for the continuous shift of critical doping level $x_{c}$ of spiral-to-AFM transition to its zero temperature value $x_{c}=0$. There is a gradual shift in the critical electron concentration, corresponding to the phase transition between spiral and AFM magnetic orders, towards half-filling as the temperature decreases. The obtained order and incommensurability parameters coincide with the results presented in Ref. BonettiThesis, of the paper.
.3 Calculation of spin susceptibility and spin current correlation functions
From Eq. (2) of the paper, it follows that the spatial spin stiffness can be directly obtained as $\chi_{\rm{op},\omega_{n}}={m^{2}}/({\bar{\chi}^{yy}_{\mathbf{Q},\omega_{n}}\omega_{n}^{2}})$. For the computation of the nonuniform dynamical spin susceptibility in the local coordinate frame, $\bar{\chi}^{\alpha,\beta}_{\mathbf{q}}$, we employ the random phase approximation (RPA), following the expression
$\bar{\chi}_{\mathbf{q}}=(\hat{I}-\bar{\chi}^{(0)}_{\mathbf{q}}\hat{U})^{-1}\bar{\chi}^{(0)}_{\mathbf{q}},$ (A10)
where all quantities are treated as matrices in the spin space, indexed by $\alpha,\beta=x,y,z,0$. Here, $\hat{U}$ represents the inter-electron interaction matrix
$\hat{U}=\left(\begin{matrix}0&0&0&-U\\ 0&U&0&0\\ 0&0&U&0\\ -U&0&0&0\end{matrix}\right),$ (A11)
while $\bar{\chi}^{(0)}_{\mathbf{q}}$ denotes the bare susceptibility bubble, defined as
$\chi^{(0)\;\alpha\beta}_{q,\omega_{n}}=-\frac{T}{4}\sum\limits_{k}\mathrm{Tr}[\sigma^{\alpha}G_{k}\sigma^{\beta}G_{k+q}]$ (A12)
with $\sigma^{\alpha}_{\sigma,\sigma^{\prime}}$ being Pauli matrices, with spin-projection indices $\sigma,\sigma^{\prime}=\uparrow,\downarrow$ and
$G_{\mathbf{k},\nu}^{\sigma\sigma^{\prime}}=\begin{pmatrix}i\nu+\mu-Un_{\downarrow}-\epsilon^{+}_{\mathbf{k}}&{\rm i}\epsilon^{-}_{\mathbf{k}}\\ -{\rm i}\epsilon^{-}_{\mathbf{k}}&i\nu+\mu-Un_{\uparrow}-\epsilon^{+}_{\mathbf{k}}\end{pmatrix}^{-1}$ (A13)
[…]
The susceptibility $\bar{\chi}_{\mathbf{q}}$ obtained from Eq. (A10) serves as a key quantity for evaluating the dynamical spin stiffness $\chi_{\text{op},\omega_{n}}$ and analyzing the stability of mean-field solutions against the emergence of inhomogeneous phases by examining the spectrum of $\bar{\chi}_{\mathbf{q}}$ throughout the entire Brillouin zone.
The spatial spin stiffness of the out-of-plane $\rho_{\text{op},\mu}$ and in-plane $\rho_{\text{ip},\mu}$ modes with $\mu=x,y$ are related to the full gauge kernel function $M^{xx}_{\mathbf{q},\omega_{n}}$ in the mean-field approximation can be expressed as OurSecond ; BonettiThesis ; Bonetti ; BonettiMetzner
$\rho_{\text{op},n}$ $=2\left({K}_{0,nn}^{(0),xx}+{K}_{0,nn}^{d}\right),$ (A14)
$\rho_{\text{ip},n}$ $={K}_{0,nn}^{(0),yy}+{K}_{0,nn}^{d},$ (A15)
where $K^{d}_{\mu\nu}$ accounts for the diamagnetic contribution, and ${K}_{0,nn}^{(0),xx}$ is a bubble contribution to the correlation function $\langle\langle\hat{J}_{\mathbf{q},{\mu}}^{\alpha}|\hat{J}_{-\mathbf{q},{\nu}}^{\beta}\rangle\rangle$ of spin currents $\hat{J}^{\alpha}_{\mathbf{q},\mu}$, $\alpha=x,y,z,0$ ($\alpha=0$ corresponds to the usual charge current operator), $\mu=x,y$. In $\rho_{in,n}$ we have also neglected the contribution $\Delta\rho_{n}$ discussed previously in Ref. OurSecond, , originating from mixing of different susceptibility components, as it was found to be numerically small. Similarly to the equation (A12) one can express ${K}_{0,nn}^{(0),xx}$ and $K^{d}_{0,nn}$ in terms of single-particle Green functions
${K}_{0,nn}^{(0),xx}$ $=\sum\limits_{s=\pm}\left[\bar{K}^{(0),xx}-is(\bar{K}^{(0),xz}-\bar{K}^{(0),zx})+\bar{K}^{(0),zz}\right]_{q,nn,ss},$ (A16)
$K_{q,\mu\nu}^{(0),yy}$ $=\bar{K}_{q,\mu\nu,++}^{(0),yy}+\bar{K}_{q,\mu\nu,--}^{(0),00}+\bar{K}_{q,\mu\nu,+-}^{(0),y0}+\bar{K}_{q,\mu\nu,-+}^{(0),0y},$ (A17)
$\bar{K}^{0;\alpha\beta}_{q;\mu\rho,ss^{\prime}}$ $=T\sum_{k}T^{\mu,s}_{{\mathbf{k},\nu},{\mathbf{q}}}\mathrm{Tr}\left[\sigma^{\alpha}G_{k}\sigma^{\beta}G_{k+q}\right]T^{\rho,s^{\prime}}_{{\mathbf{k}},{\mathbf{q}}},$ (A18)
where lower indices of all terms in the local coordinate frame are the same for all contributions and was taken out of the square brackets. Current vertices are defined as
$T^{\mu\pm}_{{\mathbf{k}},{\mathbf{q}}}$ $=(T^{\mu}_{{\mathbf{k}}-{\mathbf{Q}}/2,{\mathbf{q}}}\pm T^{\mu}_{{\mathbf{k}}+{\mathbf{Q}}/2,{\mathbf{q}}})/{2}\,\,\,$ $(\alpha,\beta=0,y),$
$T^{\mu\pm}_{{\mathbf{k}},{\mathbf{q}}}$ $=T^{\mu}_{{\mathbf{k}}+s{\mathbf{Q}}/2,{\mathbf{q}}}\,\,\,$ $(\alpha,\beta=x,z),$ (A19)
$T^{\mu}_{{\mathbf{k}},{\mathbf{q}}}=({t_{\mathbf{k}}^{\mu}+t^{\mu}_{\mathbf{k}+\mathbf{q}}})/2$ with velocity components $t^{m}_{\mathbf{k}}={\partial}\epsilon_{\mathbf{k}}/{\partial k^{m}}$.
Diamagnetic contribution can also be expressed as $K_{\mu\nu}^{d}=\bar{K}_{\mu\nu,+}^{d,00}+\bar{K}_{\mu\nu,-}^{d,yy}$, where
$\bar{K}_{\mu\nu,s}^{d,ab}=-\delta_{ab}(1-\delta_{\mu 0})(1-\delta_{\nu 0})\sum_{k}T^{\mu\nu,s}_{\mathbf{k}}\mathrm{Tr}\left[\sigma^{a}G_{k}\right].$ (A20)
.4 Spatial spin stiffnesses
For completeness, we show the dependence of spatial spin stiffnesses calculated via equations (A14) and (A15) on electron concentration, see Fig. S4. The temperature dependence of spatial spin stiffness is different for the cases of out-of-plane and in-plane responses. As temperature decreases the dependence of the in-plane spatial stiffness $\rho_{2,xx/yy}$ becomes sharper in the vicinity of half-filling. With $T\to 0$ it converges at all finite doping levels to the value nearly independent of electron concentration and develops a jump at the half-filling.
Figure S3: The dependence of spatial spin stiffness on electron concentration $n$ for various temperatures $T=\frac{1}{\beta}$ for the out-of-plane (left) and in-plane (right) excitations.
.5 Frequency dependence of the in-plane temporal spin stiffness
The frequency dependencies of the in-plane temporal spin stiffness (A10) for two electron cocentrations are presented in Fig. S4. Similarly to the out-of-plane susceptibility, discussed in the main text, it decays with the increase of frequency. At low temperatures the frequency spacing becomes finer, enabling accurate extrapolation to obtain the static limit. We note that while despite strong frequency dependence, $\chi_{\rm{op},\omega_{n}}$ is continuous at zero frequency, i.e. its zero-frequency value coincides with the considered static limit, the in-plane mode, being related to zero momentum transfer in the local coordinate frame, it however shows a jump at zeroth Matsubara frequency. The magnitude of the jump is higher at higher doping levels. Paradoxically, absence of the jump in the out-of-plane component yields its faster decay with frequency, since in that case the jump is replaced by smooth, yet rapidly decreased, function.
Figure S4: Dynamical temporal spin stiffness for in-plane mode as a function of bosonic frequency $\omega_{n}$ for various temperatures and electron concentrations.
.6 The details of the comparison of the spin gap in different approaches at various temperatures
Figure S5: The comparison of different approximations for spin gap determined from the Eq (4) of the main text at various temperatures. The dependence of energy gap logarithm $\log_{10}\Delta$ on electron concentration $n$ is computed with inclusion of only zeroth bosonic frequency $\omega_{n}=0$, all frequencies but with temporal spin stiffness approximated by its static limit $\lim_{\omega_{n}\to 0}\chi_{\rm{op},\omega_{n}}$ and the calculation with dynamic temporal spin stiffness $\chi_{\rm{op},\omega_{n}}$ obtained as a function of bosonic frequency $\omega_{n}$.
In Fig. S5 we present the dependence of the magnetic energy gap $\Delta$ (in logarithmic scale) on electron concentration $n$ for various temperatures. The results are obtained from equation (3) of the paper using three different approaches: (i) including only the zero-frequency contribution $\omega_{n}=0$, (ii) approximating the dynamic temporal spin stiffness $\chi_{\text{op},\omega_{n}}$ by its static limit $\lim\limits_{\omega_{n}\to 0}\chi_{\text{op},\omega_{n}}$, and (iii) using the exact values of $\chi_{\text{op},\omega_{n}}$ at all bosonic Matsubara frequencies $\omega_{n}$. It can be seen that both approaches (i) and (ii) results in exponential decrease of the magnetic energy gap with lowering temperature, although inclusion of finite bosonic frequencies increases $\Delta$ by several orders of magnitude. Full inclusion of temporal spin stiffness frequency dependence results in a completely different trend, that is described in the main text: only for states in a certain region of electron concentrations $n$ around $n\approx 0.9$ does the magnetic energy gap significantly decrease with lowering temperature while approaching a certain limit outside of this stability region.
[…]
unsrt